# LCLs beyond bounded degrees - Locally Finite Labelings

### v1

@@ Abstract
    We focus on the polynomial regime ($\Theta(n^{1/k}\mid k\in\mathbb{N})$) and show that whether polynomial gap results persist in the unbounded-degree setting crucially depends on how LCLs are generalized beyond bounded degrees. We first demonstrate that if one allows LCLs to be defined using infinitely many local configurations, then the polynomial gaps disappear entirely: for every real exponent $0<r\leq 1$, there exists a locally checkable problem on trees with deterministic LOCAL complexity $\Theta(n^{r})$.
    Rather than stopping at this negative result, we identify a natural class of problems for which polynomial gap results can still be recovered. We introduce Locally Finite Labelings (LFLs), which formalize the intuition that every node must fall into one of finitely many local cases, even in the presence of unbounded degrees.
@@
    Moreover, which case applies—and the corresponding value of kk—can be determined solely from the description of $\Pi$.
@@ 1 Introduction
These results give rise to what is often called the complexity landscape of LCLs on bounded-degree trees, illustrated in Figure 1. The existence of large forbidden regions in this landscape has both conceptual and algorithmic consequences: it provides a near-complete classification of local problems and allows often streamlines algorithm design. Most improtantly these results often are the result of some deeper understanding of why the complexity landscape is the way it is.
@@
Open Question How does the complexity landscape change, when we consider trees of unbounded degrees? Can we obtain similar gap results as in the bounded degree setting?
@@
Some evidence already suggests that unbounded degrees significantly alter the picture. Prior work has shown that, even on trees, the complexity of certain local problems depends explicitly on the maximum degree $\Delta$ (e.g., [3]), and that this dependence can lead to new asymptotic behaviors not present in the bounded-degree setting. In particular, by choosing $\Delta$ as a function of $n$, one can obtain deterministic complexities that do not fit into the classical tree landscape. This indicates that allowing unbounded degrees introduces genuinely new phenomena, even without cycles.
In the study of LCLs, it is customary to distinguish between the lower regime, consisting of problems solvable in $O(\log n)$ rounds or less, and the polynomial regime, consisting of problems with complexity $\omega(\log n)$. On bounded-degree trees, the polynomial regime is extremely sparse, containing only the classes $\Theta(n^{1/k})$. By contrast, on bounded-degree general graphs, this regime becomes dense and admits a wide range of intermediate complexities.
While prior work has investigated how the complexity of specific LCLs ,that make sense in the unbounded degree settting, depends on $\Delta$, most of it has focused on problems lying in the lower regime. In this paper, we initiate a systematic study of the polynomial regime for trees of unbounded degree. Our goal is to understand which aspects of the tree landscape survive once the bounded-degree assumption is removed, and which do not.
@@
A fundamental obstacle arises immediately: LCLs, as traditionally defined, do not make sense on unbounded-degree graphs, because they rely on a finite set of local configurations. Thus, before asking meaningful complexity questions, we must first decide how to generalize the notion of an LCL itself.
Open Question. What is a suitable generalization of LCLs to unbounded-degree graphs that allows one to study distributed complexity as a function of both $n$ and $\Delta$?
Several generalizations have been proposed in prior work. For example, the node-edge checkable framework underlying round elimination describes correctness via constraints on half-edge labels and node multisets, and can be extended to unbounded degrees by specifying constraints for every possible degree (see, e.g., [12]). Another approach appears in the study of binary labeling problems, where correctness depends only on adjacent edges and problems are restricted to a structurally simple class [19].
@@ 1.1 Our Contributions
A common feature of the above generalizations is that they allow an unbounded number of local configurations. Our first result shows that if such generalizations are permitted, then no meaningful polynomial gap results can exist: the complexity landscape becomes arbitrarily dense.
We do so by defining a family of problems that are locally checkable and that admit arbitrary polynomial complexities. Here, locally $3$-checkable refers to the fact that correctness can be verified by verifying ever nodes 3-hop neghborhood. The solution is then globally correct, if it is locally correct at every node.
@@
For any real $0<r\leq 1$, there exists a $3$-checkable problem using only labels $\{0,1\}$ with distributed complexity $\Theta(n^{r})$ in unbounded-degree trees.
This family of problems is defined by specifying an infinite number of $3$-hop neighborhoods that are allowed, much in the same way as LCLs, but no longer requiring the set of allowed configurations to be finite.
Motivated by this observation, we introduce Locally Finite Labelings (LFLs), which formalize the idea that every node must fall into one of finitely many local cases. As in LCLs, an LFL is specified by a finite set of labeled $r$-hop configurations. However, unlike LCLs, the $r$-hop neighborhood of a node $v$ is no longer required to be identical (up to isomorphisms) to a configuration. Instead, we mark each edge as either optional or required. Requried edges must appear exactly once in $v$s neighborhood, while optional edges may apppear arbitrarily often or not at all.
@@
Figure 2: These two configurations encode the problem of finding an MIS. Optional edges are illustrated using dashed lines and required edges are solid lines. Any node that outputs 0 must have at least one neighbor that outputs 1 and can have arbitrarily many more. On the other side any node that outputs 1 cannot have another neighbor that outputs 1, but arbitrarily many neighbors that output 0. Therefore this correctly encodes the problem of finding an MIS.
LFLs generalize LCLs to unbounded-degree graphs while preserving a finiteness condition that rules out the constructions from Theorem 1. Up to some reasonable extent, LFLs are as general as possible, while staying within the limits imposed by Theorem 1. Excitingly, we are able to show that LFLs adhere to the same gap results as LCLs in the unbounded degree setting.
@@
On the other hand, LFLs together with Theorem 2 demonstrate how much expressive power can still be allowed while preserving strong gap phenomena. In this sense, LFLs sit exactly at the boundary identified by Theorem 1. Taken together, these results characterise when and how polynomial gap results emerge in trees.
We do not view LFLs as the unique or correct generalisation of LCLs to unbounded-degree graphs — indeed, they deliberately exclude natural problems such as $(\Delta+1)$-coloring. Rather, LFLs provide the right abstraction for understanding polynomial-time gap results on trees. As such, they form an important step toward a principled theory of distributed complexity on unbounded-degree trees, and we hope that this inspires new ways to tackle the systematic study of the lower regime.
@@ 1.2 Our Contributions in more Detail
We begin by illustrating why the choice of generalisation of LCLs to unbounded-degree graphs is crucial.
The natural way to generalise LCLs
the reason why LCLs cannot encode problems in unbounded degree graphs, is because we specify each possible allowed neighborhood. The problem arises, because this set of allowed neighborhoods is required to be finite, so we could just drop this restriction. Instead we could allow a set of allowed neighborhoods to be specified for each possible degree $\Delta$. Indeed this was done in e.g.[12]. We show that this generalisation is already too expressive to lead to gap results.
A striking feature of locally checkable problems is that purely local constraints can enforce global structure. In the unbounded-degree setting, however, this phenomenon can easily lead to degenerate behaviour. We demonstrate this with a simple construction.
Consider an isolated path $P=(v_{1},\dots,v_{L})$. Attach exactly $i$ leaf nodes to $v_{i}$. Thus, the degree along the backbone of the path increases monotonically. Define a problem $\Pi$ in which a node outputs $1$ if and only if it is part of such a decorated path, and outputs 0 otherwise. This property can be enforced locally: a node with $i$ attached leaves outputs $1$ if and only if it has a neighbor that outputs 11 and has exactly i1i-1 attached leaves. Note that this constitutes an infinite number of configuraitions (one for each $i\in\mathbb{N}$).
Despite being locally checkable, this definition enforces a strong global constraint. In an $n$-node tree, such a path can have length at most $\Theta(\sqrt{n})$, since the total number of nodes grows quadratically along the path. Consequently, any distributed algorithm must explore neighborhoods of radius $\Theta(\sqrt{n})$ to determine whether a node lies on such a path, yielding distributed complexity $\Theta(\sqrt{n})$. Exploring larger neighborhoods provides no additional benefit.
This idea extends naturally. Instead of attaching $i$ leaves to $v_{i}$, we may attach $i^{\alpha}$ leaves for some $\alpha\geq 1$. This controls the maximum feasible path length and hence the complexity of the problem. Linear growth ($\alpha=1$) yields complexity $\Theta(\sqrt{n})$, while larger values of $\alpha$ give rise to problems with complexity $\Theta(n^{1/(\alpha+1)})$. By varying $\alpha$, we obtain a dense family of problems spanning a wide range of polynomial complexities.
We give the details of the above idea in Section 3. This approach is limited to obtaining only complexities in $O(\sqrt{n})$, we give a more general constuction that yiels problems with arbitrary polynomial complexity and any real exponent in Appendix B.
We defer this material since the above construction already illustrates the core issue: allowing unrestricted local constraints in the unbounded-degree setting makes it possible to encode artificial problems with finely tuned polynomial complexities, thereby eliminating the possibility of polynomial gap results.
@@
Based on the insights above, we attempt to define a generalization of LCLs to unbounded degrees. When coming up with this generalization, we had the following goals in mind:
@@
    We want to keep the spirit of having a finite amount of configurations that describes the problem. This solves the problem above. As a result the description of the problem should be independent of $\Delta$.
@@
    We want to be able to have a very expressive generalization, that captures as many problems of interest as possible.
@@
    We want to have a formalism that is not too abstract, so we can more easily reason about it.
  5. 5.
    The formalism should have a meaningful connection to the already established results about both the bounded degree case of LCLs, as well as complexities in $\Delta$.
To describe in more detail what we want to say with Goal 4 is the following. Consider the notion of Locally verifiable labelings (LVLs). A locally verifiable labeling is defined by giving a constant round LOCAL algorithm that verifies the validity of the labeling. Since it seems like whatever definition of locally correct one can feasably come up with, it should be verifiable by some mechanical procedure. This is very good at satisfying Goal 3, but, just having a black box algorithm to work with, it becomes hard to prove any interesting results. A very good example of this is the classification of LCLs in bounded degree rooted trees [6]. In that work, the authors define multiple properties of an LCL, that can be computed solely based on the description of the problem. The complexity of any given LCL can then be determined by the properties of that LCL. It is unclear how to achieve something like that using just a black box algorithm.
One note about Goal 3, while we want LFLs to encompas as many problems as possible, defining a class of problems that is as large as possible, while still being able to prove gap results. We have ommited some possible extensions to LFLs in the spirit of simplicity. For example we could have included constraints of the form ”A node must have the same number of red neigbhors as blue neighbors”. We believe that gap results also hold for problems that encode costraints like this ontop of our usual LFL constraints. But the formalism necessary to encode all such constraints would have bloated all of the definitions significantly and obfuscated a lot of the insights into the gap results.
An LFL $\Pi=(\Sigma_{\mathrm{in}},\Sigma_{\mathrm{out}},r,\mathcal{C})$ consists of an input alphabet, an output alphabet, a checking radius $r\in\mathbb{N}$, and a finite set of configurations $\mathcal{C}$.
@@
As an additional example beyond MIS, we show how $3$-coloring can be expressed as an LFL.
3-coloring
The problem of 3-coloring can now can be written up in a more intuitive sense, namely that $3COL=(\left\{\right\},\left\{R,G,B\right\},1,\mathcal{C})$ with $\mathcal{C}$ containing three configurations depicted in Figure 3.
Figure 3: These three configurations suffice to encode the problem of computing a 3-coloring in unbounded degree graphs. The arrows identify the center nodes and the deges are dashed to identify them as optional edges (there are no required edges for this problem).
@@
The second major problem is that we want to iterate all trees of a constant height [^2] Intuitively a constant height is sufficient, because there is only a finite number of of equivalence classes., this is no longer possible if the degrees are unbounded, as there are already infinitey many trees of height 2. We are tasked with defining a finite procedure that analyses an infinite number of possible trees.
@@
The first step is to move from the standard neighborhood-based definition of LFLs to a node–edge checkable variant. Conceptually, this corresponds to reducing the checking radius from an arbitrary constant $r$ to a “radius” of $1/2$, where correctness is verified solely by inspecting individual nodes and edges. This allows us to define the equivalence class on subgraphs based on the edges connecting the subgraphs to the remaining graph. Since all considered subgraphs have at most two such edges, and the enpoint of such an edge is in one of finitely many cases, this recovers our finite equivalence classes.
@@
We show that, nevertheless, for trees (and more generally for graphs of sufficiently large girth), every LFL can be encoded as a node–edge checkable LFL and vice versa, with only constant overhead. The key technical ingredient is a radius-reduction argument that allows us to focus on well-behaved local fragments of configurations. Repeating this reduction yields an equivalent LFL of radius one, for which the classical node–edge encoding can be adapted.
@@
For every LFL $\Pi=(\Sigma_{\text{in}},\Sigma_{\text{out}},r,\mathcal{C})$ in trees, there exists a node edge checkable LFL $\Pi^{\prime}=(\Sigma_{\text{in}}^{\prime},\Sigma_{\text{out}}^{\prime},\mathcal{C}_{V},\mathcal{C}_{E})$, such that:
@@
Besides being essential for our gap results, this connection makes LFLs compatible with tools such as round elimination, showing that we satisfy Goal 5 and opening up possibilities for future work.
@@
Overall, our contribution is not to introduce fundamentally new techniques, but to identify precisely how far the polynomial gap results can be pushed beyond the bounded-degree setting. We show that these gaps persist exactly up to the point where the problem definition enforces that every node falls into one of finitely many local cases. According to Theorem 1, allowing more expressive power immediately destroys the gaps, while restricting to this finiteness condition suffices to recover them even in unbounded-degree trees. In this sense, LFLs capture the maximal natural class of problems for which the polynomial gap phenomena continue to hold.

### v2

@@ Abstract
    We focus on the polynomial regime ($\Theta(n^{1/k}\mid k\in\mathbb{N})$) and show that whether polynomial gap results persist in the unbounded-degree setting crucially depends on how LCLs are generalized beyond bounded degrees.
    A complex construction in [10] shows that the polynomial gaps also vanish for LCLs on unbounded degree trees. Rather than stopping at this negative result, we give a much simpler set of problems that already contradicts the existence of any polynomial gaps. The insight obtained from this cleaner construction is that for gap results to exist, we cannot allow problems to distinguish infinitely many cases.
    This guides us to a natural class of problems for which polynomial gap results can still be recovered. We introduce Locally Finite Labelings (LFLs), which formalize the intuition that every node must fall into one of finitely many local cases, even in the presence of unbounded degrees.
@@ 1 Introduction
These results give rise to what is often called the complexity landscape of LCLs on bounded-degree trees, illustrated in Figure 1. The existence of large forbidden regions is what shapes the landscape: it provides a near-complete classification of local problems and, most importantly, these results often unveil deeper understanding of why the complexity landscape is the way it is.
@@
Some evidence already suggests that unbounded degrees significantly alter the picture. Prior work has shown that, even on trees, the complexity of certain local problems depends explicitly on the maximum degree $\Delta$ (e.g., [3]), and that this dependence can lead to new asymptotic behaviors not present in the bounded-degree setting. In particular, by choosing $\Delta$ as a function of $n$, one can obtain deterministic complexities that do not fit into the classical bounded degree tree landscape. This indicates that allowing unbounded degrees introduces genuinely new phenomena, even without cycles.
Unfortunately, this direction also leads to a negative result. Being slightly imprecise, Bousquet, Feuilloley, and Pierron [10] show that on unbounded-degree trees any complexity $f(n)\in O(n/\log n)$ can occur [^1] This is a bit imprecise, please refer to [10] for the precise statement. for what they call generalized LCLs.
However, the two negative results in bounded degree general graphs and unbounded degree trees differ in an important way. The construction for general graphs in [8] yielding arbitrary complexities for bounded degree general graphs allows a meaningful interpretation, namely that we can abuse cycles to define easier versions of problems in a structured way.
In contrast the result of [10] for the unbounded degree tree setting, offers limited to no insight into why such generalized LCLs can behave so erratically. Their work is primarily geared toward locally verifying global graph properties, such as the diameter, via a framework of local checkers. The corresponding LCL result is obtained through a reduction into this framework rather than through an explicit construction of LCLs. As a consequence, it is unclear which aspects are responsible for the collapse of structure in the complexity landscape of generalized LCLs. Do we only run into trouble, because we allow esoteric constructions in our generalisation, or is this a problem fundamental to the unbounded-degree setting?
Open Question What causes the vanishing of gap results for LCLs in unbounded degree trees?
@@
When considering LCLs in the unbounded degree setting, a fundamental obstacle arises immediately: LCLs, as traditionally defined, do not make sense on unbounded-degree graphs, because they rely on a finite set of local configurations. Thus, before asking meaningful complexity questions, we must first decide how to generalize the notion of an LCL itself. We believe that by carefully examining how exactly we define our class of problems we can get to the bottom of this issue of gap results.
To get a feel for what is important for a generalisation to the unbounded degree case, we give a clean construction of a family of locally verifiable problems for unbounded degree trees that have complexity $\Theta(n^{r})$, for any real number $0<r\leq 1$. This essentially recovers the result from [10]. Since our construction directly deals with local problems, without involving complex machinery, it offers some insight into what causes these degenerate constructions.
On some high level our construction just does the following: Define some very large number of neighborhoods such that each one can be next to each other, then order them arbitrarily by assigning them numbers 1,2,… The problem now simply consists of checking whether these neighborhoods appear in the correct order, starting with the first. By requiring nodes to only count up to a certain threshold, we can obtain any polynomial complexity we desire.
We think the main issue is the fact that our problem definition is able to make an infinite number of case distinctions, which leads us to the following question.
Open Question. Is there a suitable (and natural) generalization of LCLs to unbounded-degree graphs for which a rich complexity landscape exists?
We want to retain the essence of LCLs, that is we still just give a set of allowed configurations, for some fixed radius $r$. When the degree is bounded, the number of possible configurations is finite, so we can just list the ones that we allow and arrive at the definition of LCLs. However, in the unbounded degree setting enumerating all possibilities necessarily means moving away from a finite description. In fact, writing down just a single configuration for each possible degree would result in an infinite number of such configurations.
In [10] the authors define their generalized-LCLs in exactly this way. They just say that all possible labeled neighborhoods are classified into those that are allowed and those which are not allowed. Obviously this suffers exactly from the problem outlined above.
Some other generalizations have been proposed in prior work. For example, the node-edge checkable framework, underlying the powerful round elimination technique, describes correctness via constraints on half-edge labels, and can be extended to unbounded degrees by specifying constraints for every possible degree [12]. This formalism can be viewed as a generalisation, where nodes and edges have to only consider their radius 0.5 neighborhood to verify correctness. This formulation is just as strong as the one above and so again suffers from the same problem [^2] Altough the node-edge formalism would also require an infinite number of outputlabels to be as expressive as generalized-LCLs..
Another approach appears in the study of binary labeling problems, where correctness depends only on adjacent edges and problems are restricted to a structurally simple class [19].
@@ 1.1 Our Contributions
A common feature of the above generalizations is that they allow an unbounded number of local configurations.
Our first result shows that if such generalizations are permitted, then no meaningful polynomial gap results can exist: the complexity landscape becomes arbitrarily dense.
We do so by defining a family of problems that are locally checkable and that admit arbitrary polynomial complexities. Here, locally $4$-checkable refers to the fact that correctness can be verified by verifying every nodes 4-hop neighborhood. The solution is then globally correct, if it is locally correct at every node.
@@
For any real $0<r\leq 1$, there exists a $4$-checkable problem using only labels $\{0,1\}$ with distributed complexity $\Theta(n^{r})$ in unbounded-degree trees.
This family of problems is defined by specifying an infinite number of $4$-hop neighborhoods that are allowed, much in the same way as LCLs, but no longer requiring the set of allowed configurations to be finite.
Motivated by this observation, we introduce Locally Finite Labelings (LFLs), which formalize the idea that every node must fall into one of finitely many local cases. As in LCLs, an LFL is specified by a finite set of labeled $r$-hop configurations. However, unlike LCLs, the $r$-hop neighborhood of a node $v$ is no longer required to be identical (up to isomorphisms) to a configuration. Instead, we mark each edge as either optional or required. Requried edges must appear exactly once in $v$’s neighborhood, while optional edges may apppear arbitrarily often or not at all.
Figure 2: These two configurations encode the problem of finding an MIS. Optional edges are illustrated using dashed lines and required edges are solid lines. Any node that outputs 1 cannot have another neighbor that outputs 1, but arbitrarily many neighbors that output 0. On the other side any node that outputs 0 must have at least one neighbor that outputs 1 and can have arbitrarily many more. Therefore this correctly encodes the problem of finding an MIS.
@@
LFLs generalize LCLs to unbounded-degree graphs while preserving a finiteness condition that rules out the constructions from Theorem 1. Up to some reasonable extent, LFLs are as general as possible, while staying within the limits imposed by Theorem 1. Excitingly, we are able to show that LFLs adhere to the same polynomial gap results as LCLs in the unbounded degree setting.
@@
On the other hand, LFLs together with Theorem 2 demonstrate how much expressive power can still be allowed while preserving strong gap phenomena. In this sense, LFLs sit at the boundary identified by Theorem 1. Taken together, these results aim to characterise when and how polynomial gap results emerge in trees.
We do not view LFLs as the unique or correct generalisation of LCLs to unbounded-degree graphs — indeed, they deliberately exclude natural problems such as $(\Delta+1)$-colouring. Rather, LFLs provide the right abstraction for understanding polynomial-time gap results on trees. As such, they form an important step toward a principled theory of distributed complexity on unbounded-degree trees, and we hope that this inspires new ways to tackle the systematic study of the lower $O(\log n)$ regime.
@@ 1.2 Our Contributions in more Detail
A striking feature of locally checkable problems is that purely local constraints can enforce global structure. The combinatorial explosion introduced by unbounded degrees can easily lead to degenerate behaviour. We demonstrate this with a simple, insightful construction.
Let $\mathcal{T}_{x}$ denote the set of all rooted trees on $x\in\mathbb{N}$ nodes and height $2$. For each $x$, fix an arbitrary enumeration $(T_{i})_{i\in[|\mathcal{T}_{x}|]}$ of the trees in $\mathcal{T}_{x}$.
Consider an isolated path $P=(v_{1},\dots,v_{L})$ and attach to each node $v_{i}$ the tree $T_{i}$, with $v_{i}$ serving as the root of $T_{i}$. We refer to such a path as an $x$-gadget path. Both topology and the correct order of trees can be enforced by local constraints.
We turn this construction into a labeling problem by requiring nodes that belong to an $x$-gadget path to output 11, and all other nodes to output 0. The problem is defined so that a $x$-gadget path is valid only if it starts with $T_{1}$. Consequently, the node at the opposite end of the path can output $1$ if and only if the entire prefix $(T_{1},T_{2},\dots)$ is present. This forces any distributed algorithm to spend time linear in the length of the path in order to determine the correct output.
By requiring that only the first $L(x)$ nodes along the path output 11, we can control the maximum length of any gadget path in an instance with $n$ nodes. Choosing $L(x)$ appropriately then suffices to prove our theorem. Since the main focus are the gap results, we defer the full technical details to Appendix B.
@@
Based on the insights above, we attempt to define a generalization of LCLs to unbounded degrees, having the following goals in mind:
@@
    We want to ensure that the problem definition allows only for a finite amount of configurations that describes the problem.
@@
    We want to be able to have a generalization, that captures as many problems of interest as possible.
@@
    We want the formalism to feel natural and be easy to work with.
Since what we are aiming for is a generalisation of LCLs Goal 1 is only natural. To ensure that we do not run into the sort of degenerate behaviour described above, we have Goal 2. Note that one immediate consequence of this second goal is that our problem description has to be independent of $\Delta$.
With Goal 3 we ensure that results about LFLs are interesting, since they apply to a large amount of problems that we do actually care about. However, there is a fine line to be walked here: We also want the class of problems to have a clean and (somewhat) easy definition, so that we can reason about them. We capture this in Goal 4.
In the spirit of this trade-off between Goals 3 and 4, we have ommited some possible extensions to LFLs. For example we could prove the same gap results even while allowing for constraints of the form ”A node must have the same number of red neigbhors as blue neighbors”. Incoorporating these sorts of constraints into our definition would strengthen our results, but at the same time it would make the definitions harder to understand and clutter the proofs with heavy notation and case distinctions. We believe that our notion of LFLs strikes a nice balance.
With the above goals in mind, we propose the class of Locally Finite Labelings (LFLs). An LFL $\Pi=(\Sigma_{\mathrm{in}},\Sigma_{\mathrm{out}},r,\mathcal{C})$ consists of an input alphabet, an output alphabet, a checking radius $r\in\mathbb{N}$, and a finite set of configurations $\mathcal{C}$.
@@
As an additional example beyond MIS, we show how $3$-colouring can be expressed as an LFL.
3-colouring
The problem of 3-colouring can now be written up in a more intuitive sense, namely that $3COL=(\left\{\right\},\left\{R,G,B\right\},1,\mathcal{C})$ with $\mathcal{C}$ containing three configurations depicted in Figure 3.
Figure 3: These three configurations suffice to encode the problem of computing a 3-colouring in unbounded degree graphs. The arrows identify the center nodes and the edged are dashed to identify them as optional edges (there are no required edges for this problem).
@@
The second major problem is that we want to iterate all trees of a constant height [^4] Intuitively a constant height is sufficient, because there is only a finite number of of equivalence classes., this is no longer possible if the degrees are unbounded, as there are infinitey many trees of bounded height. We are tasked with defining a finite procedure that analyses an infinite number of possible trees.
@@
The first step is to move from the standard neighborhood-based definition of LFLs to a node–edge checkable variant. As it turns out, our notion of LFLs has a very natural translation into the world of node-edge checkable problems.
Conceptually, this reduction to the node-edge checkable setting corresponds to reducing the checking radius from an arbitrary constant $r$ to a “radius” of $1/2$, where correctness is verified solely by inspecting individual nodes and edges. This allows us to define the equivalence class on subgraphs based on the edges connecting the subgraphs to the remaining graph. Since all considered subgraphs have at most two such edges, and the endpoint of such an edge is in one of finitely many cases, this recovers our finite equivalence classes.
@@
We show that, nevertheless, for trees (and more generally for graphs of sufficiently large girth), every LFL can be encoded as a node–edge checkable LFL and vice versa. The key technical ingredient is a radius-reduction argument that allows us to focus on well-behaved local fragments of configurations. Repeating this reduction yields an equivalent LFL of radius one, which can then be turned node-edge checkable with a modified version of the original proof.
@@
For every LFL $\Pi=(\Sigma_{\text{in}},\Sigma_{\text{out}},r,\mathcal{C})$ in trees, there exists a node-edge checkable LFL $\Pi^{\prime}=(\Sigma_{\text{in}}^{\prime},\Sigma_{\text{out}}^{\prime},\mathcal{C}_{V},\mathcal{C}_{E})$, such that:
@@
Besides being essential for our gap results, this connection makes LFLs compatible with the powerful round elimination technique [12]. This opens up possibilities for future work, for example the $\omega(1)--o(\log^{*}n)$ gap result for bounded degree trees is based on round elimination. Does the same gap apply to LFLs? What is the dependence on $\Delta$?
@@
Overall, our contribution is not to introduce fundamentally new techniques, but to identify how far the polynomial gap results can be pushed beyond the bounded-degree setting. Theorem 1 illustrates how allowing for an infinite number of case distinctions leads to degenerate behaviour. Surprisingly the fact that the gap results hold for LFLs suggests that this is indeed the defining characteristic of the polynomial gap results, which hold for a large class of problems, even in the unbounded degree setting.